500 children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. What is the probability that a child, picked at random has a height greater than 120 cm?


1
Expert's answer
2022-04-13T08:15:04-0400

We have a normal distribution, "\mu=110, \sigma=6."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{120-110}{6}=1.67."


"P(X>120)=1-P(X<120)=\\=
1-P(Z<1.67)=\\
=1-0.9525=0.0475 \text{ (from z-table).}"